Solve for X and Y in the following problems using either substitution or elimination methods. Make sure you show all your work so you can get partial credit even if your final answer is wrong.
a. X + Y = 10 , 3X + Y = 12
b. 2X + 5Y = 19 , 3X + 3Y = 15
c. 4X + Y = 22 , 2X + 3Y = 16
d. 12X + Y = 174 , 8X - 2Y = 36

Answer by Alan3354(69443)

a. X + Y = 10 , 3X + Y = 12
----------------------
3x + y = 12
x + y = 10
----------- Subtract
2x = 2
x = 1
-----
y = 9
------
b. 2X + 5Y = 19 , 3X + 3Y = 15
3X + 3Y = 15 --> 2x + 2y = 10
2x + 5y = 19
2x + 2y = 10
------------ Subtract, do it like the 1st one.
--------------------------------------------
c. 4X + Y = 22 , 2X + 3Y = 16
4x + y = 22 --> y = 22 - 4x
Sub for y in the 2nd eqn
2x + 3(22 - 4x) = 16
Solve for x, then sub into either eqn to find y.
--------------------
d. 12X + Y = 174 , 8X - 2Y = 36
y = 174 - 12x
Proceed as in c.